Grade 11Algebra

If in a triangle ABC cos A=sin B/2 sin C then triangle ABC is.........(a)equilateral(b)isosceles(c)right angled(d)obtuse angled

Aniket Shriramwar
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
2sinCcosA=sinBsin(C+ A)+sin(C- A)=sinBsin(180-B)+sin(C-A)=sinBsin B+sin(C- A)=sinBsin(C-A)=0C- A=0C = A Hence isosceles triangle